OpenStudy (sasogeek):

derive an equation who's coefficients are integers and who's roots are \(\frac{1}{2} \) and \(\ 7 \) . I'm getting \(\ 2x^2-15x+7 \) but the book says \(\ 2x^2 +13x-7 \) ... who's right, who's wrong, how do u arrive at the answer?

OpenStudy (anonymous):

youre right ant the book is wrong

OpenStudy (anonymous):

Saso is always right. :)

OpenStudy (sasogeek):

i'm tired of this book and seriously want to get rid of it now but it has so many good review questions with wrong answers :@ i find more wrong answers than right answers it's so bad :@

OpenStudy (katrinakaif):

Did you confirm with Wolframalpha or does it not compute these problems?

OpenStudy (sasogeek):

i think certain questions don't even require the use of a calculator, not to even speak of wolfram

OpenStudy (katrinakaif):

I find it quite hard to believe that the book is inaccurate

OpenStudy (sasogeek):

well in ghana you can find such books and :@ i hate it!!!! believe it or not my teachers recommend it because of it's questions, not because of it's answers, sucks :@@@@@@

OpenStudy (sasogeek):

some sections are treated well, others, it's just so bad like i said, i find more wrong answers than right ones... at some sections :@@@@@@@@@@@@@@@

OpenStudy (katrinakaif):

Ah I see.

OpenStudy (anonymous):

kat, the book is inaccurate in this case, the books answer implies the roots to be -7 and 1/2

OpenStudy (katrinakaif):

I understand why it is inaccurate but I find it quite hard to believe that books trying to guide students for examinations are inaccurate.

OpenStudy (sasogeek):

do so many books have similar typos? i find this one odd

OpenStudy (katrinakaif):

Most probably

OpenStudy (anonymous):

Okay kat, tell me how I should we approach "this kinds" of problems ;)

OpenStudy (bahrom7893):

It's spelled whose btw

OpenStudy (sasogeek):

sorry my mistake lol, thanks for the heads up on that one :)

OpenStudy (sasogeek):

ffm kat might not answer u lol, it's "too early" (or too late kat) for her to solve maths

OpenStudy (katrinakaif):

I haven't been in bed..my brain feels fried.

OpenStudy (anonymous):

you dont' need brain to solve this ...

OpenStudy (katrinakaif):

What do you want?

OpenStudy (anonymous):

eyes and figures to type ..

OpenStudy (amistre64):

2x^2 +13x - 7 7*2 = 14 14 and 1 subtract to 13 (x+14/2) (x-1/2) x = -7, 1/2 ---------------------- otherwise: (x+7)(2x+1) = 2x^2 +15x - 7

OpenStudy (anonymous):

x^2 - (sum of the roots)x + product of the roots

OpenStudy (sasogeek):

roots are 7 and 1/2 => x=1/2 or x=7 =>x-7=0 or2x-1=0 hence (2x-1)(x-7)=0 and 2x^2-15x+7=0 hence expression with 1/2 and 7 as roots is 2x^2-15x+7